OpenStudy (anonymous):

x^-n(x^2m)^3 / x^2m + n I get: x^-n(x^2m)^2 / x^n aaand stuck Answer should be: x^4m-2n

OpenStudy (jack1):

could you rephrase that in the form of a sandwich?

OpenStudy (jack1):

sorry, distracted by dinner is x^-n cubed as well?

OpenStudy (anonymous):

hmm, you're right that's not very clear, I'll rewrite it :)

OpenStudy (jack1):

x^-n(x^2m)^3 / x^2m + n is it...? \[x ^{-n} \times (x ^{2m})^{3}\]

OpenStudy (anonymous):

\[x^{-n}(x^{2m})^3 / x^{2m + n}\]

OpenStudy (anonymous):

so yes :)

OpenStudy (jack1):

so first spread out the brackets: x^-n . x^6m / x^2m + n

OpenStudy (jack1):

as you've got the same base for the numerator part of the eqn: = x^(6m - n) / x^(2m +n)

OpenStudy (jack1):

and 1/ x^(2m +n) = ( x^(2m +n) )^-1

OpenStudy (jack1):

so you spread out the brackets again: = ( x^(6m - n) ) times (( x^(2m +n) )^-1) \[x ^{6m - n} \times x ^{-2m - n}\]

OpenStudy (jack1):

and again, same base, so add the powers together = x ^( (6m - n) + (- 2m - n) ) = x^(4m - 2n)

OpenStudy (jack1):

give us a heads up if you need clarification on anything, slaters

OpenStudy (anonymous):

hmm, I'll have to go over this, thank you!

OpenStudy (anonymous):

allright, finally got it :P thanks